The measurement of a supplement of an angle 6 degrees more than twice the measurement of the angle find two angles.
Answer by stanbon(75887)

The measurement of a supplement of an angle 6 degrees more than twice the measurement of the angle find two angles.
-----
Equations:
x + y = 180
x = 2y+6
------
Substitute and solve for "y":
2y+6 + y = 180
3y = 174
y = 58 degrees
-----
x = 180-58 = 122 degrees
-------------
Cheers,
Stan H.
--------------